OpenStudy (anonymous):

if the side length and apothem of a regular polygon are doubled, how will the perimeter and area change?

OpenStudy (anonymous):

 This is a drawing of the polygon before the measurements were doubled

OpenStudy (danjs):

Do you remember the area of a polygon given the apothem and the side length?

OpenStudy (anonymous):

Hello Did you get the answer?

OpenStudy (danjs):

Area = 1/2 * Apothem * perimeter perimeter = n*(side length) n=# of sides

OpenStudy (danjs):

SInce they gave you numbers, calculate each, then double the lengths, and calculate them again, and compair

OpenStudy (danjs):

Area = 1/2 *Apothem * [n*(side length)]

OpenStudy (danjs):

if you replace Apothem with 2*Apothem, and Side Length by 2*(side length) the area will have 2*2 = 4 times the original area

OpenStudy (danjs):

If the side length is doubled, the perimeter will also double since Perimeter = n *(side length) replace with 2* side length 2*Perimeter = 2*n*side length

OpenStudy (danjs):

just put the numbers in